I am trying to help my son and struggling a little since I have not looked at this stuff in years. Please help me so that I can help him.
Solve
1) -11/12 = 3h/4
2) -7 = m/4
3) 9/14a = 3/7
4) 3/5a = -6
These are the questions he missed on a recent test and I am not able to properly explain how to solve.




Answer by edjones(4080)

In all of these, if you multiply both sides time the LCD your troubles are over to eliminate the fraction(s) your troubles are over.
.
1)
-11/12 = 3h/4 LCD=12
-11=9h
h=-11/9 divide each side by 9.
.
3)
9a/14=3/7 I assume this is the problem. 14 is LCD.
9a=6
a=6/9=2/3
.
I'll let you do the other two.
.
Ed